OpenStudy (anonymous):

If (2+i)(x+yi)=1+3i, where x and y are real numbers, write two equations connecting x and y, and solve them.

OpenStudy (experimentx):

(2x-3y)+i(2y+x) = 1 + 3i compare coff ... and solve for x and y

OpenStudy (anonymous):

How?

OpenStudy (callisto):

(2+i)(x+yi)=1+3i 2x +xi +2yi + yi^2 =1+3i (2x -y) +(x +2y)i = 1+3i So 2x -y =1 x+2y = 3

OpenStudy (callisto):

treat the real part and the imaginary part separately

OpenStudy (apoorvk):

(2+i)(x+yi) = 1+3i =>2x-y + i(x+2y) = 1+3i compare real and imaginary parts. so, 2x-y=1 and x+2y=3 solve the two linear equations, to get x and y, and enjoy!

OpenStudy (callisto):

Do you understand?

OpenStudy (anonymous):

Yes, thank you :)

OpenStudy (callisto):

Welcome :)